Four years ago, Company A distributed a dividend to its shareholders of € 0.3858 per share. Today Company A paid a dividend of 0.80 € per share. In the past, the dividend policy of the company led to an increase in dividends each year at a steady rate. It is expected that the company will continue the same dividend policy for the next three years. Subsequently, the dividend growth rate will remain constant at 8% per annum for the foreseeable future. The current stock market price is 45 €. If investors require a 12% return on investing in equity-risk stocks, would you buy that share?
Answer #1

Growth rate of dividends during last 4 years (g)= [(D0/D-4)^(1/4)]-1

Where D0 = Last Dividend per share (given as €0.80) and D-4 = Dividend 4 years ago (0.3858)

Substituting the values, growth rate= (0.80/0/3858)^(1/4)-1 = 2.073613^(1/4)-1 = 20%

Given that this growth rate will apply for the next 3 years and then will be constant at 8% thereafter, for the foreseeable future.

Current value of the stock, as per Dividend Discount Model is 29.3265 €   as follows:

D 4 D1, D2 and D3 are arrived at from DO applying the growth rates g1, g2 and g3 for the respective years. 5 P3= Price after